OpenStudy (anonymous):

May rides her bike at the same distance that leah walks.May rides her bike 10 km/h FASTER than leah walks.If it takes May 1 hour and Leah 3 hours to travel that distance how fast does each travel?

OpenStudy (ranga):

Let Mary walk at x km/h. Then Leah rides at (x + 10) km/h In 1 hour Leah would have ridden her bike (x+10) * 1 = x+10 km In 3 hours Mary would have walked x * 3 = 3x km. They travel the same distance. Therefore, 3x = x + 10 solve for x. That is the speed of Mary's walking. Then put the x value in (x+10) to calculate Leah's speed.